Please use a calculator to change 4e^(-0.88i) to rectangular form rounded to 2 decimal places. Thanks!
Answer by Edwin McCravy(20060)

Use Euler's formula:

eiq = cosq + i·sinq

4e-0.88i = 4[cos(0.88) + i·sin(0.88)] = 4cos(0.88) + 4·sin(0.88)i =

2.548604577 + 3.082955516·i

which rounds to

2.55 + 3.08i

Edwin